OpenStudy (anonymous):

Kurt says that you can take the square root of each side of an equation. Therefore he says that (x+2)^2+(y-5)^2=36 and (x+2)+(y-5)=6 are equivalent equations. Dana says the equations are not equivalent. Who is right? Explain.

OpenStudy (jdoe0001):

\(\bf (x+2)^{\color{red}{ 2}}+(y-5)^{\color{red}{ 2}}=36 \implies (x+2)(x+2)+(y-5)(y-5)=36\) expand the left side, see if the equation gives \(\bf (x+2)+(y-5)=6\)

OpenStudy (triciaal):

Dana is right to prove follow the instructions above